To have a better experience, please upgrade your IE browser.upgrade
Questo sito utilizza cookie di profilazione (propri e di terze parti) per ottimizzare la tua esperienza online e per inviarti pubblicità in linea con le tue preferenze. Continuando a utilizzare questo sito senza modificare le tue preferenze acconsenti all’uso dei cookie. Se vuoi saperne di più o negare il consenso a tutti o ad alcuni cookie clicca qui>
The website that you are visiting also provides Arabian language. Do you wish to switch language version?
يوفر موقع الويب الذي تزوره المحتوى باللغة العربية أيضًا. هل ترغب في تبديل إصدار اللغة؟
The website that you are visiting also provides Russia language Do you wish to switch language version?
Данный сайт есть в английской версии. Желаете ли Вы перейти на английскую версию?
Customer has rings consists of OSN6800 that is managed by T2000 V2R7C03SPC100 , the network has many services configured as diamond and Silver services too, While both protection and Restoration via ASON is configured.
During customer maintenance actions on his fiber R_LOS alarms opserved on many links , after the maintenance window finished services switched back to the working fiber and Loss alarms dissapear , the problem that customer still complain that every day he got services inturruption on many links from 200 ~ 600 msec , he check his fiber many times and report that services run on OSN6800 is unstable while that WDM service on BWS 320G is stable.
When checking each of above causes and the log file and history alarms on each NE below answers discovered
1- the was more than one LSP avalible for diamond services.
2- the software and hardware of all nodes OSN6800 in the network was consistent. V100R006C01SPC100 (5.51.07.33)
3- Customer confirms his fiber attenuation variation via specific test instruments and also he refer that other WDM from another vendor use the same fiber cable and has stable services.
Solution\ when checking the TE-Links on the T2000 , we find that the server can query all control links with no alarms , while it can not query the TE links , in addition it showes Major alarm on each TE Link and can not query the alarm name [ error message : Alarm query fails - try again later ].
this can be shown as attached Fig-3-.
It was concluded that LMP functions and services which should run on the T2000 NMS has got problem last time the customer fiber go through maintenance for long time.
We discover the same issue when we got inconsistent message after trying to query TE trails as shown in attached Fig-4
Thus we refer to all process (Related to ASON) that is running on the server of the T2000 , and stop them all manually , then start them again.
After that action , we succeed to query the TE Links again as shown in Fig-5 , and actually customer does not record service inturruption again.
Cause -1- Only one LSP is avalible for the 1+1 Service.
Cause -2- OMS TE Link downgraded because [ the software or hardware of the board where the optical interface of the TE link is not matched with the Sink Node].
Cause -3- Customer fiber still have problems of attenuation that is changed rapidly causing the difference between working path and protection path more than 5 dB and make instantinuous switching.
Cause -4- ASON work upnormally due to loose of control channel between two nodes of the working LSP of the TE link.